OpenStudy (anonymous):

Which of the following is a polynomial with roots – sqrt 3, sqrt 3 , and –2 ? Answer x3 – 2x2 – 3x + 6 x3 + 2x2 – 3x – 6 x3 – 3x2 – 5x + 15 x3 + 3x2 – 5x – 15

OpenStudy (anonymous):

substitute x = -2 in the four equations and see which one gives you 0. similarly do it for the other roots.

OpenStudy (anonymous):

product of roots = -6, so equation 2

OpenStudy (anonymous):

lol, that is easier.

OpenStudy (mathteacher1729):

Guyc, nice solution! :) Here is the list of useful facts about polynomials: http://home.windstream.net/okrebs/page26.html Also note that you can expand \[(x-\sqrt{3})(x+\sqrt{3})(x+2)\] Since \[(x-\sqrt{3})(x+\sqrt{3})=x^2-3\] this is now just multiplying \[(x^2-3)(x+2)=x^3+2x^2-3x-6\] :)